# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ Abstract.
The fractional diffusion equation is rigorously derived as a scaling limit from a deterministic Rayleigh gas, where particles interact via short range potentials with support of size $\varepsilon$ and the background is distributed in space $\mathds{R}^{3}$ according to a Poisson process with intensity $N$ and in velocity according some fat-tailed distribution. As an intermediate step a linear Boltzmann equation is obtained in the Boltzmann-Grad limit as $\varepsilon$ tends to zero and $N$ tends to infinity with $N\varepsilon^{2}=c$. The convergence of the empiric particle dynamics to the Boltzmann-type dynamics is shown using semigroup methods to describe probability measures on collision trees associated to physical trajectories in the case of a Rayleigh gas. The fractional diffusion equation is a hydrodynamic limit for times $t\in[0,T]$, where $T$ and inverse mean free path $c$ can both be chosen as some negative rational power $\varepsilon^{-k}$.
@@ 1. Introduction
The full derivation from micro to kinetic and to other macro dynamics with or without using kinetic equations as an intermediate step, is still an open question for many equations. The rigorous derivation of the Boltzmann equation from a system of particles was first obtained by Lanford [Lan75], in the case of hard spheres and by King [Kin75] for more general potentials using the BBGKY hierarchy. The convergence in these results is valid for short times, i.e. in a time interval $[0,T^{\prime}]$, where $T^{\prime}$ is a fraction to the free flight time. Complete proofs are given in [GSRT13, PSS14]. Illner and Pulvirenti [IP86, IP89] obtained global in time results of convergence, when the initial density is small and the positions are in $\mathds{R}^{d}$. Building on [GSRT13, PSS14] fluctuations around the Maxwellian equilibrium of the Boltzmann equation can be understood and used to derive macroscopic equations and processes. Bodineau, Gallagher and Saint-Raymond [BGSR16] obtain Brownian motion as the limit of a deterministic system of hard-spheres for arbitrary large times and for an initial distribution which is close to the equilibrium state of a tagged particle. The Ornstein-Uhlenbeck process is derived for a similar setting in [BGSR18]. Further descriptions for fluctuations around the equilibrium of the Boltzmann equation are provided in [BGSRS23a, BGSRS23b]. There has been recent, substantial progress in the parallel problem of deriving the kinetic wave equation for the interaction of waves within nonlinear Schrödinger equations [LS11, DH21, DH23b, DH23a].
@@
In linear settings, variants of the Boltzmann equation can be justified for arbitrary long times, see [Spo91] and references therein. Even in the nonlinear setting of [BGSR16] the authors use the linear Boltzmann equation as an intermediate step, to pass to the Brownian motion as the limit of a deterministic system of hard-spheres, as a corollary a diffusion equation for the macroscopic distribution is obtained. The time-scale for convergence is of the order of $(\log\log N)^{-1}$. In this paper we will consider a linear setting of a tagged particle interacting with a background in a variant of the Rayleigh gas to obtain linear Boltzmann equations and then other macroscopic behaviour. A related recent result deriving diffusive behaviour of the tagged particle in a Rayleigh gas particle model with a Maxwellian background is provided in [Fou24]. In this article, Fougéres managed to find a rate of convergence which is of the order of $\exp(-c_{\beta}|\log N|^{1-\alpha})$, $\forall\alpha>0$, with $N$ be the number of particles, for times of the order $(\log|c_{\beta}\log\varepsilon|)^{\frac{1}{2}-\alpha}$ using the BBGKY hierarchy approach. This rate improves the one of [BGSR16] which is of the order $(\log\log N)^{-1}$. In [MS24], we show, starting also from a Rayleigh gas particle system with a Maxwellian background, which takes advantage of semigroup techniques as in [MST18] to study the evolution of collision trees instead of the BBGKY hierarchy, that the rate of convergence of the distribution of the tagged particle to the solution of the linear Boltzmann equation is of the order $\varepsilon^{\alpha}$, with $\alpha\in(0,\frac{11}{52})$ for time scales that are proportional to some negative power of $\varepsilon$.
@@
The present paper provides a first result of deriving a fractional diffusion equation as the limit of a deterministic particle system again on time scales that are proportional to some negative power of $\varepsilon$. Here, we are giving a specific example of short range potential Rayleigh gas with angular cut-off and with a specific fat-tailed background, using a linear Boltzmann equation as an intermediate step.
@@
We start by describing the particle dynamics. A wide, relevant class are systems with short range potentials where the particles carry a force that affects only nearby particles, up to some distance proportional to $\varepsilon$. Our current work focuses on specific short range potential models. When the particles interact via a short range potential, then typically only two particles are interacting, we will enforce this later by our specific choice of model. Thus the analysis needs only the description of the collision of $N=2$ particles. The Hamiltonian equations of motion are given by
$\frac{dx_{i}(t)}{dt}=v_{i},\quad m_{i}\frac{dv_{i}(t)}{dt}=-\nabla U(x_{1}-x_{2}),$
where $x_{i}(t)$ is the position and $v_{i}(t)$ is the velocity of particle $i$ at time $t$, for $i=1,2$ and $t\geq 0$. Here, $\varepsilon$ is the diameter of the particles, $m_{i}$ is the mass of the $i$-th particle and we consider the potential function
$U:\mathds{R}^{3}\to\mathds{R},\ U(|x|)=\begin{cases}\frac{K}{|x|^{n-1}}-K&\text{ if }|x|\leq 1\\ 0&\text{ if }|x|>1\end{cases}$ (1.1)
to be the interaction potential for $n\in(2,5]$ and $K>0$, which is radial, non-increasing, supported in the unit ball of $\mathds{R}^{3}$ and goes to zero at $|x|=1$. The scattering event for each particle can be parameterised by the scattering angle $\Theta$ defined as the angle between ingoing and outgoing relative velocities, see Fig. 1. See [GSRT13, Cer88] for extended discussions. For technical reasons to avoid the difficulties around particles undergoing collisions that take a long time or known as grazing collisions, see Figure 2, we modify the two particle interaction to ignore the potential $U$ if $\Theta\in[\frac{\pi}{2}-\sigma,\frac{\pi}{2}]$, where $\sigma$ is some fixed cut-off. Angular cut-off are often used in kinetic PDE theory in order to avoid technical difficulties related to grazing collisions. Here we already introduced it on the particle level.
@@
Particles are assumed to be small with the size of the support of the forces parameterised by $\varepsilon$. Then the particle dynamics are given by $\nabla U_{\varepsilon}(x)=\frac{1}{\varepsilon}\nabla U(\frac{x}{\varepsilon})$, as long as $\Theta\in(0,\frac{\pi}{2}-\sigma)$. These dynamics preserve overall kinetic energy and momentum. The dynamics and its potential satisfy the conditions in [GSRT13] as long as we restrict the derived collision kernels to $\Theta\in(0,\frac{\pi}{2}-\sigma)$.
Figure 2. Grazing and non-grazing collisions. The disks illustrate the support of the interaction potentials. When particles undergo a grazing collision, the particles are moving nearly in parallel for a long time.
@@
If the background particles are of infinite relative mass to the tagged particle and the background particles are fixed, then this model is known as the Lorentz gas model. Long-term convergence was shown in [Spo78]. For recent progress about random Lorentz gas see the paper of Lutsko and Toth, [LT20]. In this work, they prove the invariance principle for a random Lorentz-gas particle in three space dimension under the Boltzmann-Grad limit and simultaneous diffusion scaling. They do this by using a coupling of the mechanical trajectory and some controls on the efficiency of this coupling. Another recent result on the Lorentz gas is [MS21] where they provide a derivation of transport processes from a Lorentz gas particle system for radial potentials with compact support in the low density limit.
A Rayleigh gas is related to the Lorentz gas, but the background particles are no longer of infinite mass. If the background particles are of equal mass with the tagged particle and the background particles interact only with the tagged particle and not with each other then this model is known as the Rayleigh gas. We consider a tagged particle with initial distribution $f_{0}$ and $g_{0}$ the distribution of the background particles, distributed via a spatial homogeneous Poisson process. We index the particle by $i\in\mathds{N}$. There is no interaction among the background particles. For simplicity we will ignore multiple collision of the tagged particle at the same time. Such events have vanishing probability in the scaling limit. To summarise, we first define the collision status $\beta_{i}(t)$ that is piece-wise constant in time for $i\in\mathds{N}$ with jumps from 00 to $1$ and $1$ to 00 if
[…]
We will ignore any initial overlaps, i.e. particles will not interact in that case. We start with data $\beta_{i}(0)=0$ for all $i\in\mathds{N}$, then the collision state are well defined for all times for almost all initial data as the only undefined case –when $|x_{i}(t)-x_{0}(t)|=\varepsilon$ for several $i$– has probability 00. Setting $\beta_{0}(t)=\max_{j\in\mathds{N}}\beta_{j}$ we can define the many particle dynamics as follows
[…]
@@
$Q^{+}[f_{t}](x,v)=\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega},$
@@
$Q^{-}[f_{t}](x,v)=f_{t}(x,v)\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}g_{0}(\bar{v})B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}.$
Here $B$ is the collision kernel which is derived from a two particle interaction and has the following form
$B(v_{1}-v_{2},\Theta)$ $=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\left[|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\right]\Theta(\sin\Theta)^{-1},$ (1.5)
for $n\in(2,5]$, large $|v_{1}-v_{2}|>0$ and $\Theta\in(0,\frac{\pi}{2}-\sigma)$. See section 3 for the derivation of this formula. Furthermore, we are making the assumption that the initial distribution of the background particles, $g_{0}:\mathds{R}^{3}\to\mathds{R}$, is radial and has the following form
@@
[…]
@@
There is a widely developed theory to derive continuum equations, such as the Navier-Stokes, Euler and Heat equations as scaling limits from the Boltzmann equation. For some review of methods and the substantial literature see e.g. [CIP94, SR09, Sle13, Gal19]. For particular relevance for us is the work [MMM11], where Mellet, Mischler and Mouhot provide a proof for the convergence of the linear Boltzmann equation to a fractional diffusion equation, where the equilibrium distribution function is a heavy-tailed distribution with infinite variance.
@@
$\begin{cases}\partial_{\tau}\rho+\kappa(-\Delta_{x})^{\frac{\gamma}{2}}\rho&=0\quad\quad\mathrm{in}\ (0,\infty)\times\mathds{R}^{3},\\ \quad\quad\quad\quad\quad\rho(0,\cdot)&=\rho_{0}\quad\quad\mathrm{in}\ \mathds{R}^{3},\end{cases}$ (1.7)
with
$\kappa=\frac{\kappa_{0}\nu_{0}}{1-\beta}\int_{\mathds{R}^{3}}\frac{w_{1}^{2}}{\nu_{0}^{2}+w_{1}^{2}}\frac{1}{|w|^{3+\gamma}}\,\mathrm{d}w.$
[…]
@@
Let $f_{0}\in L^{1}(\mathds{R}^{3}\times\mathds{R}^{3})$ the initial distribution of the tagged particle with $f_{0}(x,v)(1+|v|)\in L^{1}(\mathds{R}^{3}\times\mathds{R}^{3})$ and $g_{0}(v)(1+|v|)\in L^{1}(\mathds{R}^{3})$, where $g_{0}$ is the distribution of the background particles and let $t\in[0,T_{\varepsilon}]$, $c^{\frac{28}{29}}T_{\varepsilon}=\varepsilon^{\frac{38m}{29}-\frac{5}{29}}$, $0<m<\frac{5}{38}$. Then the distribution of the tagged particle $\hat{f}^{N}_{t}$ converges in the $L^{1}$-norm to the solution of the linear Boltzmann equation (1.4) $f_{t}$, in the Boltzmann-Grad limit $N\varepsilon^{2}=c$, for a time $T_{\varepsilon}$ which is diverging with $\varepsilon\to 0$. That is, there exists $\varepsilon_{0}>0$ such that for every $\varepsilon>0$, with $\varepsilon<\varepsilon_{0}$ such that for any $t\in[0,T_{\varepsilon}]$ the error can be estimated by
@@
The following theorem is about the derivation of a fractional diffusion equation from a Rayleigh gas particle system with the background particles being initially distributed according to a spatial homogeneous Poisson process. In the macroscopic limit, the trajectory of the tagged particle is defined by $\Xi(\tau):=x(\tilde{\beta}\tau)\in\mathds{R}^{3}$. The distribution of $\Xi(\tau)$ is given by $\hat{f}^{N}(\tilde{\beta}\tau,x,v)$, see Lemma 5.4 for the choice of $\tilde{\beta}$ and how it is related with the parameter $c$.
@@
Let $f_{0}\in L^{1}\cap L^{2}_{F^{-1}}(\mathds{R}^{3}\times\mathds{R}^{3})$ the initial distribution of the tagged particle and consider the background particles are Poisson distributed according to the density $Ng_{0}(v)\mathrm{d}x\mathrm{d}v$. Define $\gamma(q):=\frac{\alpha-\beta}{1-\beta}$, $\alpha=q-3$, $\beta<5-q$ and $4<q<\frac{4n}{n-1}$, with $n\in(2,5]$. Let $F$ be an equilibrium distribution of the linear Boltzmann equation. Let also $\rho(\tau,x)$ be the solution of the fractional diffusion equation (1.7). Then the distribution of the tagged particle $\hat{f}^{N}(\tilde{\beta}\tau,x,v)$ converges in $L^{\infty}(0,T;L^{2}_{F^{-1}}\cap L^{1}(\mathds{R}^{3}\times\mathds{R}^{3}))$-norm to $\rho(\tau,x)F(v)$, i.e. for any test function $\tilde{\varphi}\in L^{1}(0,T;L^{2}_{F^{-1}}\cap L^{\infty}(\mathds{R}^{3}\times\mathds{R}^{3}))$
$|(\hat{f}^{N}(\tilde{\beta}\tau,x,v)-\rho(\tau,x)F(v))\cdot\tilde{\varphi}(\tau,x,v)|\to 0,$
@@
In the next section we provide an overview of the main ingredients for the proofs of both theorems. In section 3 we appose some general introduction on how to define the collision kernel $B$. Furthermore we find the formula of the collision kernel $B$ for the case $n\in(2,5]$, for large relative velocities and for deviation angle $\Theta\in(0,\frac{\pi}{2}-\sigma)$, for $\sigma>0$. In section 4 we prove Theorem 1.1 by adapting and extending [MST18] with quantitative error estimates and by considering short range potential dynamics on the phase space $\mathds{R}^{3}\times\mathds{R}^{3}$. The proof of Theorem 1.2 is given in section 5. We adapt results from [MMM11] and combine them with Theorem 1.1.

### v2

@@ Abstract.
The fractional diffusion equation is rigorously derived as a scaling limit from a deterministic Rayleigh gas, where particles interact via short range potentials with support of size $\varepsilon$ and the background is distributed in space $\mathds{R}^{3}$ according to a Poisson process with intensity $N$ and in velocity according to some fat-tailed distribution. As an intermediate step a linear Boltzmann equation is obtained in the Boltzmann-Grad limit as $\varepsilon$ tends to zero and $N$ tends to infinity with $N\varepsilon^{2}=c$. The convergence of the empiric particle dynamics to the Boltzmann-type dynamics is shown using semigroup methods to describe probability measures on collision trees associated to physical trajectories in the case of a Rayleigh gas. The fractional diffusion equation is a hydrodynamic limit for times $t\in[0,T]$, where $T$ and inverse mean free path $c$ can both be chosen as some negative rational power $\varepsilon^{-k}$.
*Corresponding author: Theodora Syntaka, theodora.syntaka@bath.edu
@@ 1. Introduction
The full derivation from micro to kinetic and to other macro dynamics with or without using kinetic equations as an intermediate step, is still an open question for many equations. The rigorous derivation of the Boltzmann equation from a system of particles was first obtained by Lanford [Lan75], in the case of hard spheres and by King [Kin75] for more general potentials using the BBGKY hierarchy. The convergence in these results is valid for short times, i.e. in a time interval $0,T^{\prime}]$, where $T^{\prime}$ is a fraction to the free flight time. Complete proofs are given in [GSRT13, PSS14]. Illner and Pulvirenti [IP86, IP89] obtained global in time results of convergence, when the initial density is small and the positions are in $\mathds{R}^{d}$. There has been recent, substantial progress in the parallel problem of deriving the kinetic wave equation for the interaction of waves within nonlinear Schrödinger equations [LS11, DH21, DH23b, DH23a]. Recently, Hani, Deng and Ma announced substantial progress for the long-term derivation of the Boltzmann equation [DHM24] and then for derivation of Euler and Navier-Stokes equations [DHM25] using the Boltzmann equation as an intermediate step.
Building on [GSRT13, PSS14] fluctuations around the Maxwellian equilibrium of the Boltzmann equation can be understood and used to derive macroscopic equations and processes. Bodineau, Gallagher and Saint-Raymond [BGSR16] obtain Brownian motion as the limit of a deterministic system of hard-spheres for arbitrary large times and for an initial distribution which is close to the equilibrium state of a tagged particle. The Ornstein-Uhlenbeck process is derived for a similar setting in [BGSR18]. Further descriptions for fluctuations around the equilibrium of the Boltzmann equation are provided in [BGSRS23a, BGSRS23b].
@@
In linear settings, variants of the Boltzmann equation can be justified for arbitrary long times, see e.g. [Spo91] and references therein. Even in the nonlinear setting of [BGSR16] the authors use the linear Boltzmann equation as an intermediate step, to pass to the Brownian motion as the limit of a deterministic system of hard-spheres, as a corollary a diffusion equation for the macroscopic distribution is obtained. The time-scale for convergence is of the order of $(\log\log N)^{-1}$. In this paper we will consider a linear setting of a tagged particle interacting with a background in a variant of the Rayleigh gas to obtain linear Boltzmann equations and then other macroscopic behaviour. A related recent result deriving diffusive behaviour of the tagged particle in a Rayleigh gas particle model with a Maxwellian background is provided in [Fou24]. In this article, Fougères managed to find a rate of convergence which is of the order of $\exp(-c_{\beta}|\log N|^{1-\alpha})$, $\forall\alpha>0$, with $N$ be the number of particles, for times of the order $(\log|c_{\beta}\log\varepsilon|)^{\frac{1}{2}-\alpha}$ using the BBGKY hierarchy approach. This rate improves the one of [BGSR16] which is of the order $(\log\log N)^{-1}$. In [MS24], we show, starting also from a Rayleigh gas particle system with a Maxwellian background, which takes advantage of semigroup techniques as in [MST18] to study the evolution of collision trees instead of the BBGKY hierarchy, that the rate of convergence of the distribution of the tagged particle to the solution of the linear Boltzmann equation is of the order $\varepsilon^{\alpha}$, with $\alpha\in(0,\frac{1}{8})$ for time scales that are proportional to some negative power of $\varepsilon$.
@@
The present paper provides a first result of deriving a fractional diffusion equation as the limit of a deterministic particle system again on time scales that are proportional to some negative power of $\varepsilon$. Here, we are giving a specific example of short range potential Rayleigh gas and with a specific fat-tailed background, using a linear Boltzmann equation as an intermediate step.
@@
We start by describing the particle dynamics. A wide, relevant class are systems with short range potentials where the particles carry a force that affects only nearby particles, up to some distance proportional to $\varepsilon$. Our current work focuses on specific short range potential models. When the particles interact via a short range potential, then typically only two particles are interacting. Thus the main analysis needs only the description of the collision of $N=2$ particles. The Hamiltonian equations of motion are given by
$\frac{dx_{i}(t)}{dt}=v_{i},\quad m_{i}\frac{dv_{i}(t)}{dt}=-\nabla U_{\varepsilon}(x_{i}-x_{j}),$
where $(x_{i}(t),v_{i}(t))\in\mathds{R}_{x}^{3}\times\mathds{R}_{v}^{3}$ is the position and velocity of particle $i$ at time $t$, for $i=1,2$, where $i\neq j=1,2$, and $t\geq 0$. Here, $\varepsilon$ is the diameter of the particles, $m_{i}$ is the mass of the $i$-th particle and we consider the potential function
$U:\mathds{R}^{3}\to\mathds{R},\ U(|x|):=\bar{K}(|x|^{1-n}-1)\mathds{1}_{|x|\leq 1}$ (1.1)
to be the interaction potential for $n\in(2,5]$ and $\bar{K}>0$, which is radial, non-increasing, supported in the unit ball of $\mathds{R}^{3}$ and goes to zero at $|x|=1$. Since $\bar{K}$ is a positive constant, we can set $\bar{K}=1$ along the article for simplification. The scattering event for each particle can be parametrised by the scattering angle $\Theta$ defined as the angle between ingoing and outgoing relative velocities, see Fig. 1. See [GSRT13, Cer88, PSS14] for extended discussions.
@@
Particles are assumed to be small with the size of the support of the forces parametrised by $\varepsilon$. Then the particle dynamics are given by $\nabla U_{\varepsilon}(x)=\frac{1}{\varepsilon}\nabla U(\frac{x}{\varepsilon})$. These dynamics preserve overall kinetic energy and momentum. The dynamics and its potential satisfy the conditions in [GSRT13].
@@
If the background particles are of infinite relative mass to the tagged particle and the background particles are fixed, then this model is known as the Lorentz gas model. Long-term convergence was shown in [Spo78]. For recent progress about random Lorentz gas see the paper of Lutsko and Toth, [LT20]. In this work, the authors prove the invariance principle for a random Lorentz-gas particle in three space dimension under the Boltzmann-Grad limit and simultaneous diffusion scaling. They do this by using a coupling of the mechanical trajectory and some controls on the efficiency of this coupling. Another recent result on the Lorentz gas is by Marklof and Strömbergsson [MS24] which provides a derivation of transport processes from a deterministic Lorentz gas particle system for radial potentials with compact support in the low density limit .
A Rayleigh gas is related to the Lorentz gas, but the background particles are no longer of infinite mass. If the background particles are of equal mass with the tagged particle and the background particles interact only with the tagged particle and not with each other then this model is known as the Rayleigh gas. We consider a tagged particle with initial distribution $f_{0}$ and $g_{0}$ the distribution of the background particles, distributed via a spatial homogeneous Poisson process. We index the particles by $i\in\mathds{N}$. There is no interaction among the background particles.
In the main part of our analysis we can ignore multiple collision of the tagged particle at the same time. Such events have vanishing probability in the scaling limit. Due to the compact support of $\nabla U_{\varepsilon}$ the tagged particle will only interact with a finite number of particles with probability $1$. The equations of motions are given by
$\begin{cases}\dot{x}_{i}(t)&=v_{i}(t)\text{ for all }i\in\{0\}\cup\mathds{N}\\ \dot{v}_{0}(t)&=-\sum_{j\in\mathds{N}}\nabla U_{\varepsilon}(x_{0}(t)-x_{j}(t))\\ \dot{v}_{j}(t)&=-\nabla U_{\varepsilon}(x_{j}(t)-x_{0}(t))\text{ if }j\in\mathds{N}.\end{cases}$ (1.2)
The initial condition $(x_{0},v_{0})$ of the tagged particle 0 is chosen according to some $f_{0}\in L^{1}(\mathds{R}^{3}_{x}\times\mathds{R}^{3}_{v})$. The background particles are placed spatially in $\mathds{R}^{3}$ according to a Poisson process with intensity $N$, the velocities are chosen independently according to the probability density $g_{0}\in L^{1}(\mathds{R}^{3}_{v})$
@@
$Q^{+}[f_{t}](x,v)=\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega},$
@@
$Q^{-}[f_{t}](x,v)=f_{t}(x,v)\int_{\mathbb{S}^{2}}\int_{\mathds{R}^{3}}g_{0}(\bar{v})\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}.$
Here $\mathcal{B}$ is the collision kernel which is derived from a two particle interaction and has the following form
$\mathcal{B}(v_{1}-v_{2},\Theta)$ $=\frac{2^{\frac{2}{n-1}}}{a^{2}}\left[|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\right]\Theta(\sin\Theta)^{-1},$ (1.4)
for $n\in(2,5]$ and $|v_{1}-v_{2}|\to\infty$. Here, $a$ is a constant defined by $a:=\int_{0}^{1}\frac{\mathrm{d}w}{\sqrt{1-(1-w)^{n-1}}}$, for $n\in(3,5]$. See section 3 for the derivation of this formula. Furthermore, we are making the assumption that the initial distribution of the background particles, $g_{0}:\mathds{R}^{3}\to\mathds{R}$, is radial, continuous and has the following form
@@
where $C_{1},C_{2}$, $\bar{R}$ are positive, real constants and $q\in(4,\frac{4n}{n-1})$, $n\in(3,5]$ such that $\int_{\mathds{R}^{3}}g_{0}(v)\mathrm{d}v=1$. Notice that the solutions of the linear Boltzmann equation (1.3) conserves mass but not energy. The derivation of (1.3) is given in [MST18] for the case of hard sphere dynamics, for finite, fixed times without any error estimates. For the derivation of the linear Boltzmann equation from a Lorentz gas particle system see [Gal69]. Also in [ET17, DP99] the linear Boltzmann equation (1.3) is derived from a long range particle evolution. For variants and further details see also [Sto17, MS18] and for a related model [NWL19, DR01]. Various scaling limits can be considered for systems with long-range potentials [NVW21]. For the derivation of the linear Boltzmann equation without cut-off from a particle system where particles interact by a infinite range potential see [Ayi17].
@@
There is a widely developed theory to derive continuum equations, such as the Navier-Stokes, Euler and Heat equations as scaling limits from the Boltzmann equation. For some review of methods and the substantial literature see e.g. [CIP94, SR09, Sle13, Gal19]. For particular relevance for us is the work [MMM11], where Mellet, Mischler and Mouhot provide a proof for the convergence of the linear Boltzmann equation to a fractional diffusion equation, where the equilibrium distribution function is a heavy-tailed distribution with infinite variance. For a recent review on the hydrodynamic limit toward Fractional Diffusion equations see [Dech23]. We also refer to [CdPF23] for the derivation of the fractional Porous Medium equation as the hydrodynamic limit of a random particle system with long range interactions.
@@
$\begin{cases}\partial_{\tau}\rho+\kappa(-\Delta_{x})^{\frac{\gamma}{2}}\rho&=0\quad\quad\mathrm{in}\ (0,\infty)\times\mathds{R}^{3},\\ \quad\quad\quad\quad\quad\rho(0,\cdot)&=\rho_{0}\quad\quad\mathrm{in}\ \mathds{R}^{3}.\end{cases}$ (1.6)
The function $\rho=\rho(\tau,x)$ represents the density $\int_{\mathds{R}^{3}}f(\tau,x,v)\mathrm{d}v$. The fractional operator is defined as $(-\Delta_{x})^{\frac{\gamma}{2}}\rho:=\mathcal{F}^{-1}\big(|k|^{\gamma}\mathcal{F}(\rho)(k)\big)$, with $\mathcal{F}$ the Fourier transform in the space variable, and it is non-local operator since it is defined through the Fourier transform. The fractional operator of the order of $\frac{\gamma}{2}\in(0,1)$ generates a rotationally symmetric $\gamma$-stable Lévy process, see e.g. [Sch16] or [WGMN97]. For a reference for the state of the art of the fractional diffusion equations we refer to [EL18].
In this article, our main theorem concerns the convergence from a deterministic microscopic system to a fractional diffusion (1.6) continuum limit. In this microscopic system, the background particles are distributed in position according to a spatial homogeneous Poisson process with intensity $N$ and in velocity according to a fat-tailed distribution $g_{0}$ as in (1.5). The derivation of the fractional diffusion equation depends on the interaction of two ingredients: the decay assumptions of this distribution $g_{0}$ and the specific choice particle interaction given in (1.1).
@@
Let $f_{0}\in L^{1}(\mathds{R}^{3}\times\mathds{R}^{3})$ the initial distribution of the tagged particle with $f_{0}(x,v)(1+|v|)\in L^{1}(\mathds{R}^{3}\times\mathds{R}^{3})$ and $g_{0}(v)(1+|v|)\in L^{1}(\mathds{R}^{3})$, where $g_{0}$ is the distribution of the background particles and let $t\in[0,T_{\varepsilon}]$, $cT_{\varepsilon}=\varepsilon^{\frac{4m}{3}-\frac{2}{9}},$ $0<m<\frac{1}{6}$. Then the distribution of the tagged particle $\hat{f}^{N}_{t}$ converges in the $L^{1}$-norm to the solution of the linear Boltzmann equation (1.3) $f_{t}$, in the Boltzmann-Grad limit $N\varepsilon^{2}=c$, for a time $T_{\varepsilon}$ which is diverging with $\varepsilon\to 0$. That is, there exists $\varepsilon_{0}>0$ such that for every $\varepsilon>0$, with $\varepsilon<\varepsilon_{0}$ such that for any $t\in[0,T_{\varepsilon}]$ the error can be estimated by
@@
We compare our result with those of [GSRT13] and [MST18]. In [GSRT13], the authors rigorously derive the Boltzmann equation from a system of hard spheres or Newtonian particles interacting via a short-range potential. In [MST18], the linear Boltzmann equation is derived in the Boltzmann–Grad limit from a Rayleigh gas particle system. The primary distinction lies in the methodological approach: [GSRT13] relies on the BBGKY hierarchy, whereas both our work and [MST18] use semigroup methods for the comparison of probabilities of genuine collision histories.
A key difference also concerns the time scale for which convergence to the Boltzmann equation is established. In [GSRT13], convergence holds for short times. In contrast, [MST18] proves convergence for arbitrarily long but finite, fixed times. Our result goes further, establishing convergence for times that diverge in the Boltzmann–Grad limit.
The next theorem establishes the rigorous justification of a fractional diffusion equation from a Rayleigh gas particle system with the background particles being initially distributed according to a spatial homogeneous Poisson process. In the macroscopic limit, the trajectory of the tagged particle is defined by $\Xi(\tau):=x(\tilde{\beta}\tau)\in\mathds{R}^{3}$. The distribution of $\Xi(\tau)$ is given by $\hat{f}^{N}(\tilde{\beta}\tau,x,v)$, see Lemma 5.10 where the admissible choice of $\tilde{\beta}$ is given, which is $\tilde{\beta}=\epsilon^{1-{\gamma}}$, and the relation of this with the parameter $c$.
@@
Let $f_{0}\in L^{1}\cap L^{2}_{F^{-1}}(\mathds{R}^{3}\times\mathds{R}^{3})$ be the initial distribution of the tagged particle and consider the background particles that are Poisson distributed according to the density $Ng_{0}(v)\mathrm{d}x\mathrm{d}v$. Define $\gamma(q):=\frac{(q-4)(n-1)}{4}+1$, for $4<q<\frac{4n}{n-1}$ and $n\in(3,5]$. Let $F$ be the unique positive equilibrium distribution of the linear Boltzmann equation. Let $\rho(\tau,x)$ be the solution of the fractional diffusion equation (1.6). Then the distribution of the tagged particle $\hat{f}^{N}(\epsilon^{1-{\gamma}}\tau,x,v)$ converges in $L^{\infty}(0,T;L^{2}_{F^{-1}}\cap L^{1}(\mathds{R}^{3}\times\mathds{R}^{3}))$-norm to $\rho(\tau,x)F(v)$, i.e. for any test function $\tilde{\varphi}\in L^{1}(0,T;L^{2}_{F^{-1}}\cap L^{\infty}(\mathds{R}^{3}\times\mathds{R}^{3}))$
$\int_{0}^{T}\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|(\hat{f}^{N}(\epsilon^{1-{\gamma}}\tau,x,v)-\rho(\tau,x)F(v))\cdot\tilde{\varphi}(\tau,x,v)|\mathrm{d}x\mathrm{d}v\mathrm{d}\tau\to 0,$
@@
In the next section we provide an overview of the main ingredients for the proofs of both theorems. In section 3 we provide a general introduction on how to obtain the collision kernel $\mathcal{B}$. Furthermore we find the formula of the collision kernel $\mathcal{B}$ for the case $n\in(3,5]$, for large relative velocities. In section 4 we prove Theorem 1.1 by adapting and extending [MST18] with quantitative error estimates and by considering short range potential dynamics on the phase space $\mathds{R}^{3}\times\mathds{R}^{3}$. The proof of Theorem 1.2 is given in section 5. We adapt results from [MMM11] and combine them with Theorem 1.1.